# AED-PADA:Improving Generalizability of Adversarial Example Detection via Principal Adversarial Domain Adaptation

### v1

@@ 4.3 Analysis
In this experiment, we utilize 3 distinct feature extractions, i.e., spatial feature extraction, frequency feature extraction, and the adversarial feature enhancement (spatial + freq) in our framework. Still 7 unseen attacks are employed and the averaged results are reported in Table 5. As can be observed, our adversarial feature enhancement (spatial + freq) exhibit superior performances than both the spatial feature extraction and frequency feature extraction, which verifies the effectiveness of our work.
The compatibility with different MDA methods.
@@
Here, we employ two widely used MDA methods, MFSAN [?] and M3DA-$\beta$ [?], to demonstrate the compatibility of our framework with the existing MDA methods. Table 6 presents that the performances of employing both the MFSAN and M3DA-$\beta$ can surpass the existing SOTA adversarial detection methods. This verifies that our framework possesses excellent compatibilities with existing MDA methods. Besides, since MFSAN outperforms M3DA-$\beta$ in the majority of experiments, we select MFSAN as the default MDA component in our PADA.
5 Conclusion
@@
In this paper, we proposed a novel and effective adversarial example detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation, to improve the generalization ability of adversarial detection. Specifically, AED-PADA contains two stages, i.e., Principal Adversarial Domains Identification and Principal Adversarial Domain Adaptation. In PADI, we acquired ADs from scratch and constructed PADs as the source domains for PADA. In PADA, we constructed an adversarial feature enhancement based Multi-source Domain Adaptation, which is compatible with various existing MDA methods, to effectively leverage PADs to achieve adversarial detection. Experimental results proved the superiority of our work, compared to the state-of-the-art detection methods.

### v2

@@ 4.5. Effectiveness of the adversarial feature enhancement in PADA
In this experiment, we utilize three distinct feature extractions, i.e., spatial feature extraction, frequency feature extraction, and the adversarial feature enhancement (spatial + freq) in our framework. Still 7 test unseen attacks are employed and the averaged results are reported in Table 12. As can be observed, in both datasets, the average results of our adversarial feature enhancement(spatial + freq) exhibit superior performances than both the spatial feature extraction and frequency feature extraction. This indicates that using adversarial feature enhancement as the feature extraction method for PADA is effective. Furthermore, whether it is spatial feature extraction or frequency feature extraction, the performance of single domain feature extraction across different backbones is inconsistent, as evidenced by the fluctuated averaged detection performance across different backbones on SVHN. Our adversarial feature enhancement, which combines both spatial and frequency aspects, can more comprehensively capture adversarial perturbation signals, yielding superior adversarial detection performance.
4.6. The compatibility with different MUDA methods
@@
To demonstrate the compatibility of our framework with the existing Multi-source Unsupervised Domain Adaptation (MUDA) methods, we employ four widely used Multi-source Unsupervised Domain Adaptation (MUDA) methods, M3DA (moment-mda, ), DARN (DARN, ), MDMN (mdmn, ) and MFSAN (aaai-mda, ). All four MUDA methods have utilized the adversarial feature enhancement as the feature extraction component. Table 13 presents that the performances of employing all four MUDA methods can surpass the existing SOTA adversarial detection methods. This verifies that our framework possesses excellent compatibility with existing MUDA methods. In other words, The first exploration of MUDA methods in adversarial example detection has proven to be successful, and Principal Adversarial Domain Adaptation can effectively transfer the knowledge from PADs to the unseen target domain. Besides, since the average result of MFSAN outperforms other MUDA methods, we select MFSAN as the basic MUDA component in our PADA.
Table 13. Detection performance of our framework with different MUDA methods.
Dataset Backbone Averaged Accuracy of different MUDA methods (%)
M3DA (moment-mda, ) DARN (DARN, ) MDMN (mdmn, ) MFSAN (Ours) (aaai-mda, )
CIFAR-10 ResNet-18 94.976 93.948 93.832 94.694
CIFAR-10 VGG-16 87.819 87.624 91.463 88.864
SVHN ResNet-18 93.892 93.934 93.888 94.344
SVHN VGG-16 90.960 92.244 89.826 91.219
Average 91.912 91.938 92.252 92.280
4.7. Parameters sensitivity.
We utilize MFSAN as the basic MUDA method in the PADA stage of our framework, and $\lambda=\gamma=1.0$ control the importance of $L_{d}$ and $L_{disc}$, respectively. To study the parameters sensitivity, we sample the values of $\lambda$ and $\gamma$ from {0.5, 1.0, 2.0}, and perform the experiments under the settings of CIFAR-10 and ResNet-18.
Table 14. Sensitivity of the parameters $\lambda$ and $\gamma$ in the PADA stage.
$\lambda$ $\gamma$ Averaged Accuracy (%)
$\lambda$=0.5 $\gamma$=0.5 94.503
$\lambda$=0.5 $\gamma$=1.0 94.591
$\lambda$=0.5 $\gamma$=2.0 94.355
$\lambda$=1.0 $\gamma$=0.5 94.646
$\lambda$=1.0 $\gamma$=1.0 94.694
$\lambda$=1.0 $\gamma$=2.0 94.603
$\lambda$=2.0 $\gamma$=0.5 94.598
$\lambda$=2.0 $\gamma$=1.0 94.629
$\lambda$=2.0 $\gamma$=2.0 94.223
Table 14 indicates that our proposed method maintains high performance consistency under various parameter settings when employing MFSAN as the basic MUDA method. The best averaged accuracy is 94.694% when $\lambda=\gamma=1.0$. Variations of the parameters typically do not induce the fluctuations in the detection performance. The mean value and the standard deviation of all the averaged accuracies are 94.538% and 0.154%, respectively. This indicates that our PADA framework is effective and robust, since it shows insensitivity to different parameters and consistently offers high detection accuracy.
4.8. The computational costs
Table 15. Comparison of the time and hardware costs. Batch size for each method is 1000.
Cost Detector
LID (lid, ) MD (md, ) Steg (steg, ) SID (SID, ) SA (txt_advdetection, ) Ours
Time (ms/image) 8.13 18.05 23.85 2.16 1.21 0.60
Hardware (MB) 7458 7248 1614 1592 8564 6124
Table 15 presents the time and hardware costs when deploying the state-of-the-art adversarial example detection methods and our AED-PADA. The results indicate that our method has the minimal time expenses and moderate hardware (GPU memory) expenses under the same settings. All experiments are conducted on an NVIDIA GeForce RTX 3080Ti GPU under CIFAR-10 and ResNet-18 settings.
Furthermore, our AED-PADA is capable of deployment in real-world scenarios. Real-world scenarios demand the adversarial example detection methods not only possess the capability for real-time detection but also maintain robust detection performance in the environments where the adversarial attacks, datasets, and backbones during testing are entirely unseen. Firstly, we only train once, ready for real-world deployment without retraining for new unseen attacks. We have the shortest inference time, which is the best real-time detection performance. Secondly, we train from earlier attacks and test on more advanced attacks. Table 3 demonstrates that under this real-world-aligned setting, our AED-PADA exhibits superior generalization performance, and would maintain considerable detection capabilities against future unseen attacks. Lastly, with its strong performance across various backbones and datasets depicted in the Table 6, our proposed method performs better in the environments with unseen datasets and backbones. Consequently, our AED-PADA is well-suited for practical applications in the complex real-world environments.
5. Conclusion
@@
In this paper, we proposed a novel and effective adversarial example detection method, named Adversarial Example Detection via Principal Adversarial Domain Adaptation, to improve the generalization ability of adversarial detection. Specifically, AED-PADA contains two stages, i.e., Principal Adversarial Domains Identification (PADI) and Principal Adversarial Domain Adaptation (PADA). In PADI, we acquired ADs from scratch and constructed PADs as the source domains for PADA. In PADA, we proposed an adversarial feature enhancement based Multi-source Unsupervised Domain Adaptation framework, which is compatible with various existing MUDA methods, to effectively leverage PADs to achieve adversarial example detection. Experimental results demonstrated the superiority of our work, compared to the state-of-the-art detection methods.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China under Grant 62272020, U20B2069 and 62176253, in part by the State Key Laboratory of Complex & Critical Software Environment under Grant SKLSDE2023ZX-16, and in part by the Fundamental Research Funds for Central Universities.